# Seat Arrangement Problems under B-utility and W-utility

### v1

@@ 4 Envy-freeness (EFA)
Case 2: let $z$ be between $b$ and $c$ in $\sigma^{\prime}$. As $b$ was envious of $c$ in $\sigma$, $f_{b}(c)+f_{b}(a)<f_{b}(c)+f_{b}(d)$, so $d(b,d)<d(a,b)$. If $z$ is not to be envious of $b$ in $\sigma^{\prime}$, then $c$ is to the right of $a$ in the $1$-dimensional space, as otherwise $f_{z}(a)>f_{z}(c)$. Similarly, $b$ is to the right of $d$, or $z$ would envy $c$ under $\sigma^{\prime}$. If $b$ is to not be envious of $z$, then $d(a,b)\leq d(b,c)$. Similarly, $d(c,d)\leq d(b,c)$, or $c$ would envy $z$. If, in the $1$-dimensional space, $b$ is to the left of $a$, or between $a$ and $c$, then, as $b$ is to the right of $d$, $d(c,d)>d(b,c)$, a contradiction. Thus, $b$ is to the right of $c$, and because $c$ is to the right of $a$, $d(a,b)>d(b,c)$. However, we stated before that $d(a,b)\leq d(b,c)$, or else $b$ is envious of $z$ in $\sigma^{\prime}$, a contradiction. Thus, regardless of the positions of the agents in the $1$-dimensional space, under $\sigma^{\prime}$ at least one agent will envy one of its neighbours.
@@
Case 3: assume that $z$ is placed between $a$ and $b$, and that this causes $b$ to stop being envious of $c$. Hence $f_{b}(z)+f_{b}(c)\geq f_{b}(c)+f_{b}(d)$. As $z$ is the rightmost agent, $b$ must be to the right of $d$ in the $1$-dimensional space. We also have that $a$ is to the left of $d$, as $f_{b}(z)\geq f_{b}(d)$ implies that $b$ is to the right of $d$, and because $b$ used to envy $c$, we have that $f_{b}(a)<f_{b}(d)$. As $a$ cannot be to the right of $b$, because in that case $f_{b}(a)>f_{b}(z)\geq f_{b}(d)>f_{b}(a)$, it must be to the left of $d$. $c$ is to the right of $a$, too, or else $f_{c}(a)+f_{c}(b)>f_{c}(b)+f_{c}(d)$, which would mean that in $\sigma$ $c$ envies $b$, which would be covered by the first case. Thus, we have that $f_{z}(a)+f_{z}(b)<f_{z}(b)+f_{z}(c)$, so $z$ is envious of $b$.
@@
Let $I=(P,F,G,S)$ be an instance of Seat Arrangement with S-utility, $1$-dimensional preferences with unique positions, and a seat graph containing at least one path component with at least three vertices, or a cycle component with at least four. Then no envy-free assignment exists. The result on paths extends to positive preferences in general.
@@
We first prove the result for paths. Let $a,b,c\in P$ be agents in a Seat Arrangement instance with positive preferences, and let $\sigma$ be an arrangement that assigns $a$ to the end vertex of a path component of the seat graph, $b$ to the neighbour of the vertex assigned to $a$, and $c$ to the other neighbour of the vertex assigned to $b$. As preferences are strictly positive, and the utility type is S-utility, $u_{\sigma^{a\leftrightarrow b}}(a)=f_{a}(b)+f_{a}(c)>f_{a}(b)=u_{\sigma}(a)$. Hence, $a$ envies $b$.
Regarding the result for cycles, let $I=(P,F,G,S)$ be an instance of Seat Arrangement, such that preferences are derived from a $1$-dimensional space, and $G$ contains a cycle component with four vertices or more. Let $\sigma$ be any arrangement and let $\sigma^{\prime}$ be the arrangement $\sigma$ restricted to the cycle component. Then, by Lemma 2, $\sigma^{\prime}$ is not envy-free, and hence $\sigma$ is not envy-free either. As $\sigma$ was arbitrary, $I$ does not have an envy-free arrangement.∎
We now consider the case of B-utility. It turns out that, even with $1$-dimensional and strict preferences, it is NP-complete to determine the existence of an envy-free arrangement. In order to prove this statement, we will reduce from the Bin Packing problem [14]. In an instance of Bin Packing, we are given a set $X$ of items, such that each item $x_{i}\in X$ has a size $s_{i}\in\mathbb{N}$, a capacity $B\in\mathbb{N}$ for each bin, and a number of bins $K$. The set of all item sizes is $S$. The goal is to decide whether $X$ can be partitioned into at most $K$ different subsets such that the sum of the sizes of the elements in each subset is not greater than $B$.
@@
Let $I=(P,F,G,B)$ be a Seat Arrangement instance with B-utility, $1$-dimensional strict preferences with unique positions, and a seat graph in which every connected component is a path with the same number of vertices. Then it is NP-complete to decide whether there exists an envy-free arrangement.
@@
Let $(X,S,B,K)$ be an instance of Bin Packing, where $I$ is the set of elements, $B$ the bin capacity, and $K$ the target number of bins. As a further restriction, the minimum size for each item is $2$; we can ensure this by multiplying every $s_{i}\in S$, as well as the bin capacity $B$, by $2$. The bins in the Bin Packing will be labelled $B_{1},\dots,B_{K}$. Construct the Seat Arrangement instance $I=(P,F,G,B)$ as follows. Take an item $x_{i}\in X$ and create $s_{i}$ agents, $p_{i}^{j}\in P$ ($1\leq j\leq s_{i}$). All of these agents belong to subset $P_{i}\subset P$. Lay them out in the real line starting with $p_{i}^{1}$, which will be placed at 00. Then $p_{i}^{2}$ will be placed a distance of exactly $1$ to the right of $p_{i}^{1}$, $p_{i}^{3}$ will be placed a distance of $1+\varepsilon$ to the right of $p_{i}^{2}$, where $\varepsilon$ is an arbitrarily small but strictly positive number, $p_{i}^{4}$ will be placed a distance of $1+2\cdot\varepsilon$ to the right of $p_{i}^{3}$, and so on, until $p_{i}^{s_{i}}$, which will be placed a distance of $1+(s_{i}-2)\cdot\varepsilon$ to the right of $p_{i}^{s_{i}-1}$. After completing the placement of all agents associated with item $x_{i}$, we then proceed with the placement of item $x_{i+1}$. Its first agent, $p_{i+1}^{1}\in P$, will be placed a distance of $2$ to the right of $p_{i}^{s_{i}}$, and we repeat the process with $p_{i+1}^{2}$ and so on. The seat graph $G$ consist of $K$ paths with $B$ vertices each, labelled $P_{B}^{1}$ to $P_{B}^{K}$.
Now suppose that there is a packing of the items into $k$ bins, for $k\leq K$. Without loss of generality assume that bins $B_{1},\dots,B_{k}$ are used. For each $\ell$ ($1\leq\ell\leq k$), suppose that the $\ell$th bin contains items $x_{j_{1}},\dots,x_{j_{r}}$ for some $r\geq 1$, where $j_{1}<j_{2}<\dots<j_{r}$. Place the agents in $P_{j_{1}},P_{j_{2}},\dots,P_{j_{r}}$ onto the path $P_{B}^{\ell}$ in left-to-right order that respects the placement of these agents on the real line as constructed above. As every item $x_{i}$ is packed into a bin, all agents from $P_{i}$ are assigned a vertex in $G$ such that they feel no envy towards any other agent.
Conversely, suppose that there is an envy-free arrangement of the agents onto $G$. Let $k$ ($1\leq k\leq K$) be given and suppose that path $P_{B}^{k}$ contains an agent $p_{i}^{r}$. Then by the envy-free property, $P_{B}^{k}$ contains every agent in $P_{i}$. It follows that the packing obtained by placing item $x_{i}$ into bin $B_{k}$ whenever the family $P_{i}$ is placed on path $P_{B}^{k}$ is a legal packing of all items in to at most $K$ bins. Therefore, as Bin Packing is NP-complete, so is Seat Arrangement with B-utility, $1$-dimensional strict preferences, and a seat graph in which every connected component is a path with the same number of vertices. ∎
5 Maximum (MWA) and maximin (MUA) assignments
We now show our results on MWA and MUA. For any Seat Arrangement instance that has utility type S-utility, MWA is NP-hard even with symmetric and binary preferences [4], and, if Spanning Subgraph Isomorphism of regular graphs is NP-hard, so is MUA, even with symmetric and binary preferences [4]. We improve the result for MUA and conclude that the problem is definitely NP-hard. We also show that, with S-utility, both MWA and MUA remain NP-hard even with symmetric and strict preferences. Finally, we extend these results to B-utility and W-utility.
Theorem 9.
For either B-utility, S-utility, or W-utility, MWA and MUA are NP-hard, even if preferences are binary and symmetric, or symmetric and strict, and the seat graph $G$ is planar and has bounded degree.
Proof.
@@ 5 Maximum (MWA) and maximin (MUA) assignments
Start first with B-utility. We reuse the reduction from Theorem 3. Recall that, in that reduction, from Partition Into Triangles to EFA, there exists an envy-free arrangement $\sigma$ if and only every agent has utility $1$ under $\sigma$, which happens if and only if there exists a partition into triangles in the Partition Into Triangles instance $I$. Thus, in our Seat Arrangement instance $I=(P,F,H,B)$, with $n$ agents that have binary and symmetric preferences, there exists an arrangement $\sigma$ in which $\sum_{p\in P}u_{\sigma}(p)=n$ if and only if there exists an arrangement $\sigma$ where every agent has utility at least $1$. Furthermore, there exists an arrangement $\sigma$ where every agent has utility at least $1$ if and only if there exists a partition into triangles in the PIT instance $G$. Thus, both MWA and MUA are NP-hard. These results hold if we arbitrarily break ties in preferences, in a manner similar to that of Theorem 4.
@@
In the case of W-utility, we use the same reduction. Like in the previous case, if preferences are binary and symmetric there exists an arrangement $\sigma$ in which $\sum_{p\in P}u_{\sigma}(p)=n$ if and only if there exists an arrangement $\sigma$ where every agent has utility at least $1$, which can only happen if there is a partition into triangles in $I$. Thus, an arrangement with total utility $n$ exists if and only if a partition into triangles exists, proving that MWA is NP-hard, while an arrangement with minimum utility exists if and only if an arrangement into triangles exists, proving that MUA is NP-hard. By arbitrarily breaking ties in preferences, we prove that both results hold if preferences are symmetric and strict.
@@
For the case of S-utility, binary and symmetric preferences, and MWA, our goal is to find an arrangement $\sigma$ in which $\sum_{p\in P}u_{\sigma}(p)=2|E(H)|$, where $E(H)$ is the set of edges of the seat graph. This can happen if and only if the score of every edge is $1$ (remember that edges are counted twice in the previous sum). In such an arrangement, if we use a B-utility function every agent has utility $1$, which implies that there exists a partition in triangles in $G$. As the converse direction also holds, MWA is NP-hard. The result holds if preferences are symmetric and strict, proven by arbitrarily breaking ties in preferences.
@@
What remains is the case with S-utility, binary and symmetric preferences (or symmetric and strict), and MUA. Our goal is to find an arrangement $\sigma$ where every agent gets utility $1$ or more. By Claim $2$ of Theorem 3, with minor modifications, this can happen if and only if there is a partition into triangles in $G$. Like in the previous cases, we can show that the result holds for symmetric and strict preferences by arbitrarily breaking ties. ∎
@@
From an instance of Bin Packing, with item set $X$, item size set $S$, bin capacity $B$, and number of bins $K$ we create a Seat Arrangement instance $I=(P,F,G,B)$ using the reduction from Theorem 8. In general, we have that, if preferences are $1$-dimensional, for any two agents $p,q$, $f_{p}(q)=f_{q}(p)=D-d(p,q)+1$, where $D$ is the maximum distance between agents, and $d(p,q)$ is the distance from $p$ to $q$ in the $1$-dimensional instance. Given how preferences are constructed, for agents $p_{i}^{l}$ and $p_{i}^{l+1}$, which are created from the same element $i\in I$, $f_{p_{i}^{l}}(p_{i}^{l+1})=f_{p_{i}^{l+1}}(p_{i}^{l})>D-(1+n\cdot\varepsilon)+1=D-n\cdot\varepsilon$, where $n=|P|$, the number of agents in $I$, and $\varepsilon$ is the arbitrarily small but strictly positive constant used in the proof of Theorem 8. For any other agents $p$ and $q$, $f_{p}(q)=f_{q}(p)\leq D-2+1=D-1$. Thus, there exists a matching in which every agent has utility at least $D-n\cdot\varepsilon$ if and only if agents can be arranged in an envy-free arrangement like the one described in Theorem 8, which happens if an only if the elements in $X$ can be packed into bins. Hence, MUA is NP-hard.
As for MWA, there exists an arrangement $\sigma$ in which the sum of all utilities $\sum_{p\in P}u_{\sigma}(p)$ is at least $n\cdot(D-n\cdot\varepsilon)$ if an only if the utility of every agent is at least $D-n\cdot\varepsilon$, which in turn happens if and only if the elements in $I$ can be packed into bins, proving that MWA is NP-hard. ∎
@@ 6 Conclusion
We have introduced two new types of utilities to Seat Arrangement, namely B-utility and W-utility, giving us a more complete picture of the algorithmic behaviour of the problem for different utility functions. In particular, we have seen that, in general, the complexity of a given variant of Seat Arrangement does not change between S-utility and both B-utility and W-utility. However, some problems that are computationally hard with S-utility become tractable with B-utility. One example is finding an exchange-stable arrangement when preferences are symmetric. Additionally, we have studied a new type of agent preference, derived from a one dimensional space, which we call $1$-dimensional preferences. In this case, if the seat graph contains a cycle or path component, and the utility type is S-utility, no envy-free arrangement can exist. Meanwhile, if preferences are $1$-dimensional and the utility type is $B$-utility, EFA, MWA, and MUA are NP-hard.

### v2

@@ 4 Envy-freeness (EFA)
Case 2: let $z$ be between $q$ and $r$ in $\sigma^{\prime}$. The positions of the agents in $G$ are shown is Figure 6. As $q$ was envious of $r$ in $\sigma$, $f_{q}(r)+f_{q}(p)<f_{q}(r)+f_{q}(s)$, so $d(q,s)<d(p,q)$. If $z$ is not to be envious of $q$ in $\sigma^{\prime}$, then $r$ is to the right of $p$ in the $1$-dimensional space, as otherwise $f_{z}(p)>f_{z}(r)$. Similarly, $q$ is to the right of $s$, or $z$ would envy $r$ under $\sigma^{\prime}$. If $q$ is to not be envious of $z$, then $d(p,q)\leq d(q,r)$. Similarly, $d(r,s)\leq d(q,r)$, or $r$ would envy $z$. If, in the $1$-dimensional space, $q$ is to the left of $p$, or between $p$ and $r$, then, as $q$ is to the right of $s$, $d(r,s)>d(q,r)$, a contradiction. Thus, $q$ is to the right of $r$, and because $r$ is to the right of $p$, $d(p,q)>d(q,r)$. However, we stated before that $d(p,q)\leq d(q,r)$, or else $q$ is envious of $z$ in $\sigma^{\prime}$, a contradiction. Thus, regardless of the positions of the agents in the $1$-dimensional space, under $\sigma^{\prime}$ at least one agent will envy one of its neighbours.
ppqqzzrrss
Figure 6: Positions of the agents in part of the seat graph $G$. Case 22.
@@
Case 3: assume that $z$ is placed between $p$ and $q$ in $\sigma^{\prime}$ The positions of the agents in $G$ are shown is Figure 5. As $q$ does not envy $r$ anymore, $f_{q}(z)+f_{q}(r)\geq f_{q}(r)+f_{q}(s)$, so $f_{q}(z)\geq f_{q}(s)$. As $z$ is the rightmost agent, $q$ must be to the right of $s$ in the $1$-dimensional space. We now consider the position of $p$. First, as $q$ used to envy $r$, we have that $f_{q}(p)<f_{q}(s)$. Thus, $p$ cannot be to the right of $q$, because in that case $f_{q}(p)>f_{q}(z)\geq f_{q}(s)>f_{q}(p)$. Similarly, she cannot be between $q$ and $s$, because that implies that $f_{q}(p)<f_{q}(s)$, a contradiction. Hence, $p$ must be to the left of $s$. $r$ is to the right of $p$, too, or else $f_{r}(p)+f_{r}(q)>f_{r}(q)+f_{r}(s)$, which would mean that in $\sigma$ $r$ envies $q$, which by symmetry with the first case implies that $\sigma^{\prime}$ is not envy-free. Thus, we have that $f_{z}(p)+f_{z}(q)<f_{z}(q)+f_{z}(r)$, so $z$ is envious of $b$.
ppzzqqrrss
Figure 7: Positions of the agents in part of the seat graph $G$. Case 33.
@@
Let $I=(P,F,G,S)$ be an instance of seat arrangement with S-utility, $1$-dimensional preferences with unique positions, and a seat graph containing at least one path component with at least three vertices, or a cycle component with at least four. Then no envy-free assignment exists.
@@
We first prove the result for paths. Let $a,b,c\in P$ be agents in a seat arrangement instance with $1$-dimensional preferences with unique positions, and let $\sigma$ be an arrangement that assigns $a$ to the end vertex of a path component of the seat graph, $b$ to the neighbour of the vertex assigned to $a$, and $c$ to the other neighbour of the vertex assigned to $b$. As preferences are strictly positive, and the utility type is S-utility, $u_{\sigma^{a\leftrightarrow b}}(a)=f_{a}(b)+f_{a}(c)>f_{a}(b)=u_{\sigma}(a)$. Hence, $a$ envies $b$.
Regarding the result for cycles, let $I=(P,F,G,S)$ be an instance of seat arrangement, such that preferences are derived from a $1$-dimensional space, and $G$ contains a cycle component with four vertices or more. Let $\sigma$ be any arrangement. Then, by Lemma 3, an agent assigned to the cycle component with $4$ or more vertices envies another agent assigned to the same cycle, so $\sigma$ is not envy-free. As $\sigma$ was arbitrary, $I$ does not have an envy-free arrangement.∎
In the proof of the result for path graphs in Theorem 8 we only used the fact that $1$-dimensional preferences are positive. Thus, we get the following corollary.
Corollary 1.
Let $I=(P,F,G,S)$ be an instance of seat arrangement with S-utility, positive preferences, and a seat graph containing at least one path component with at least three vertices. Then $I$ has no envy-free arrangement.
5 Maximum (MWA) and maximin (MUA) assignments
We now show our results on MWA and MUA. For any seat arrangement instance that has utility type S-utility, MWA is NP-hard even with symmetric and binary preferences [bodlaender_hedonic_2020], and, if Spanning Subgraph Isomorphism of regular graphs is NP-hard, so is MUA, even with symmetric and binary preferences [bodlaender_hedonic_2020]. We improve the result for MUA and conclude that the problem is definitely NP-hard. We also show that, with S-utility, both MWA and MUA remain NP-hard even with symmetric and strict preferences. Finally, we extend these results to B-utility and W-utility.
@@ 5 Maximum (MWA) and maximin (MUA) assignments
For either B-utility, S-utility, or W-utility, MWA and MUA are NP-hard, even if preferences are binary and symmetric, or symmetric and strict, and the seat graph is planar and has bounded degree.
@@
Start first with B-utility. We reuse the reduction from Theorem 3. Recall that, in that reduction, from Partition Into Triangles to EFA, there exists an envy-free arrangement $\sigma$ if and only every agent has utility $1$ under $\sigma$, which happens if and only if there exists a partition into triangles in the Partition Into Triangles instance $G$. Thus, in our seat arrangement instance $I=(P,F,H,B)$, with $n$ agents that have binary and symmetric preferences, there exists an arrangement $\sigma$ in which $\sum_{p\in P}u_{\sigma}(p)=n$ if and only if there exists an arrangement $\sigma$ where every agent has utility at least $1$. Furthermore, there exists an arrangement $\sigma$ where every agent has utility at least $1$ if and only if there exists a partition into triangles in the PIT instance $G$. Thus, both MWA and MUA are NP-hard.
@@
If preferences are symmetric and strict, we can modify the valuations from Theorem 3 by using arbitrarily small, positive and pairwise distinct $\varepsilon_{i}$, in a manner similar to that of Theorem 5. With these new preferences, it is still the case that there exists an arrangement $\sigma$ in which $\sum_{p\in P}u_{\sigma}(p)=n$ if and only if there exists an arrangement $\sigma$ where every agent has utility at least $1$, exists if and only if there exists a partition into triangles in $G$. Therefore, under B-utility, both MWA and MUA are NP-hard even if preferences are symmetric and strict.
In the case of W-utility, we use the same reduction. Like in the previous case, if preferences are binary and symmetric there exists an arrangement $\sigma$ in which $\sum_{p\in P}u_{\sigma}(p)=n$ if and only if there exists an arrangement $\sigma$ where every agent has utility at least $1$, which can only happen if there is a partition into triangles in $G$. Thus, an arrangement with total utility $n$ exists if and only if a partition into triangles exists, proving that MWA is NP-hard, while an arrangement with minimum utility $1$ exists if and only if an arrangement into triangles exists, proving that MUA is NP-hard. By using arbitrarily small, positive and pairwise distinct $\varepsilon_{i}$ to arbitrarily break ties in preferences, we prove that both results hold if preferences are symmetric and strict.
@@
For the case of S-utility, binary and symmetric preferences, and MWA, our goal is to find an arrangement $\sigma$ in which $\sum_{p\in P}u_{\sigma}(p)=2|E(H)|$, where $E(H)$ is the set of edges of the seat graph $H$. This can happen if and only if the score of every edge is $1$ (recall that edges are counted twice in the previous sum). The proof of Theorem 3 shows that a partition into triangles in $G$ implies an arrangement in which the score of every edge is $1$, and that an envy-free arrangement under B-utility implies that the score of every edge is $1$, which in turns implies the existence of a partition into triangles in $G$. Thus, there exists an arrangement $\sigma$ in which $\sum_{p\in P}u_{\sigma}(p)=2|E(H)|$ if and only if there exists a partition in triangles in $G$, so MWA is NP-hard. The result holds if preferences are symmetric and strict, proven by arbitrarily breaking ties in preferences.
@@
What remains is the case with S-utility, binary and symmetric preferences (or symmetric and strict), and MUA. Our goal is to find an arrangement $\sigma$ where every agent gets utility $1$ or more, which happens if every agent gives a valuation of $1$ to at least one neighbour. By Claim $2$ of Theorem 3 this can happen if and only if there is a partition into triangles in $G$. Like in the previous cases, we can show that the result holds for symmetric and strict preferences by arbitrarily breaking ties. ∎
@@
From an instance of Bin Packing, with item set $X$, item size set $S$, bin capacity $C$, and number of bins $K$ we create a seat arrangement instance $I=(P,F,G,B)$ using the reduction from Theorem 4. In general, we have that, if preferences are $1$-dimensional, for any two agents $p,q$, $f_{p}(q)=f_{q}(p)=D-d(p,q)+1$, where $D$ is the maximum distance between agents, and $d(p,q)$ is the distance from $p$ to $q$ in the $1$-dimensional instance. Given how preferences are constructed, for agents $p_{i}^{l}$ and $p_{i}^{l+1}$, which are created from the same element $i\in I$, $f_{p_{i}^{l}}(p_{i}^{l+1})=f_{p_{i}^{l+1}}(p_{i}^{l})>D-(1+n\cdot\varepsilon)+1=D-n\cdot\varepsilon$, where $n=|P|$, the number of agents in $I$, and $\varepsilon$ is the arbitrarily small but strictly positive constant used in the proof of Theorem 4. For any other agents $p$ and $q$, $f_{p}(q)=f_{q}(p)\leq D-2+1=D-1$. Thus, there exists a matching in which every agent has utility at least $D-n\cdot\varepsilon$ if and only if agents can be arranged in an envy-free arrangement like the one described in Theorem 4, which happens if an only if the elements in $X$ can be packed into bins. Hence, MUA is NP-hard.
As for MWA, there exists an arrangement $\sigma$ in which the sum of all utilities $\sum_{p\in P}u_{\sigma}(p)$ is at least $n\cdot(D-n\cdot\varepsilon)$ if an only if the utility of every agent is at least $D-n\cdot\varepsilon$, which in turn happens if and only if the elements in $X$ can be packed into bins, proving that MWA is NP-hard. ∎
@@ 6 Conclusion
We have introduced two new types of utilities to seat arrangement, namely B-utility and W-utility, giving us a more complete picture of the algorithmic behaviour of the problem for different utility functions. In particular, we have seen that, in general, the complexity of a given variant of seat arrangement does not change between S-utility and both B-utility and W-utility. However, the computational complexity of some problems differs between the different utility types. One example is finding an exchange-stable arrangement when preferences are symmetric, because with S-utility it is PLS-complete, but with B-utility it is solvable in polynomial time. Additionally, we have studied a type of agent preference, derived from a one dimensional space, which we call $1$-dimensional preferences. In this case, if the seat graph contains a cycle or path component, and the utility type is S-utility, no envy-free arrangement can exist. Meanwhile, if preferences are $1$-dimensional and the utility type is $B$-utility, EFA, MWA, and MUA are NP-hard.